### v1

# MatchFixAgent: Language-Agnostic Autonomous Repository-Level Code Translation Validation and Repair

Abstract.
Code translation transforms source code from one programming language (PL) to another. Validating the functional equivalence of translation and repairing, if necessary, are critical steps in code translation. Existing automated validation and repair approaches struggle to generalize to many PLs due to high engineering overhead, and they rely on existing and often inadequate test suites, which results in false claims of equivalence and ineffective translation repair. To bridge this gap, we develop MatchFixAgent, a large language model (LLM)-based, PL-agnostic framework for equivalence validation and repair of translations. MatchFixAgent features a multi-agent architecture that divides equivalence validation into several sub-tasks to ensure thorough and consistent semantic analysis of the translation. Then it feeds this analysis to test agent to write and execute tests. Upon observing a test failure, the repair agent attempts to fix the translation bug. The final (in)equivalence decision is made by the verdict agent, considering semantic analyses and test execution results.
We compare MatchFixAgent’s validation and repair results with four repository-level code translation techniques. We use 2,2192{,}219 translation pairs (each consisting of a source function and its translation) from their artifacts, which cover $6$ PL pairs, and are collected from $24$ GitHub projects totaling over $900K$ lines of code. Our results demonstrate that MatchFixAgent produces (in)equivalence verdicts for $99.2\%$ of translation pairs, with the same equivalence validation result as prior work on $72.8\%$ of them. When MatchFixAgent ’s result disagrees with prior work, we find that $60.7\%$ of the time MatchFixAgent ’s result is actually correct. In addition, we show that MatchFixAgent can repair $50.6\%$ of inequivalent translation, compared to prior work’s $18.5\%$. This demonstrates that MatchFixAgent is far more adaptable to many PL pairs (with a small overhead of 1,6501{,}650 lines of code) than prior work, while producing highly accurate validation results.
Program Analysis, Neuro-Symbolic Code Translation Validation and Repair, LLM Agent
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Code translation, the process of converting source code from one programming language (PL) to another, is a cornerstone of software modernization efforts that enhance performance, maintainability, and reliability (Khan et al., 2022; Jain and Chana, 2015; Jamshidi et al., 2013; Khadka et al., 2014). Translation validation and repair are integral steps in code translation for determining functional equivalence and patch generation for incorrect translations. However, performing validation and repair manually—particularly in large codebases—can be tedious, time-consuming, and error-prone, especially when complex code structures and dependencies are involved (Hora et al., 2015; Kula et al., 2018; Wang et al., 2020). Prior work on repository-level code translation defines value and type equivalences and translations to compare source and target implementations over pairs of concrete inputs. The inputs either come from existing tests from the source project (Shetty et al., 2024; Zhang et al., 2025; Wang et al., 2025a; Ibrahimzada et al., 2025) or differential fuzzing (Yang et al., 2024c; Eniser et al., 2024). Despite notable advancements in validation and repair of repository-level code translation, existing techniques are hampered by the following limitations:
  1. (1)
    Difficulty Generalizing to Many PL Pairs. While the fundamental ideas of current validation approaches extend to many PL pairs, their actual implementations typically support just one language pair. This is because supporting language interoperability between a pair of PLs requires a large engineering effort, as evidenced by the size of these tools [^1] The implementation of notable recent translation and validation techniques are AlphaTrans ($10859$ LoC), Oxidizer ($19052$ LoC), and Skel ($3843$ LoC).. Given the quadratic number of PL pairs, language interoperability techniques are extremely challenging to scale to many PL pairs.
  2. (2)
    Unknown Test Requirements. Current translation validation approaches require a set of valid inputs to validate the input-output equivalence between source functions and their translation. They generate these inputs by either executing available source tests (Wang et al., 2025a; Zhang et al., 2025; Ibrahimzada et al., 2025; Shetty et al., 2024) or fuzzing the source project (Eniser et al., 2024; Yang et al., 2024c) Unit tests are often incomplete, missing important inputs, and resulting in false claims of equivalence. Fuzzing techniques suffer from generating invalid inputs, also resulting in false claims of inequivalence, and in general, failing to reach deep into the code or create complex objects in the context of real-world projects (Klees et al., 2018; Godefroid et al., 2008).
  3. (3)
    Ineffective Translation Repair. Recent studies have shown that a more rigorous validation can reveal more translation bugs (Pan et al., 2024). Hence, advancement in translation validation should be accompanied by effective repair strategies. Existing techniques, however, either require the user to fix incorrect translations manually (Wang et al., 2025a) or use feedback-driven re-prompting strategies (Zhang et al., 2025; Ibrahimzada et al., 2025) that are barely effective in repository-level code translation due to long call-chain dependencies in real-world projects (Ibrahimzada et al., 2025).
Prior work consists of high-effort implementations that can deliver inconsistent results and only apply to one of a quadratic number of language pairs. Large Language Models (LLMs) have recently been successful at same-language equivalence validation (Wei et al., 2025a; Maveli et al., 2025), so replacing cross-language equivalence implementations with LLM decisions is a logical next step. While there is a risk of incorrect results, the baseline mechanical approaches already display low accuracy.
Towards this end, we propose MatchFixAgent, a language-agnostic approach to automate validation and repair of repository-level code translation (§3.1). MatchFixAgent combines the generative power of LLMs with several approximate code semantic analyses, i.e., analysis of control- and data-flow paths, library APIs, exception handling, and (formal or informal) specification (§3.2). These semantic analyses are then fed to a test generator & repair agent to generate tests for assessing functional equivalence, and repair the translation in the case of failing tests (§3.3). The final equivalence decision will be made by the verdict agent, considering the approximate semantic analyses and test execution results (§3.4). MatchFixAgent is very lightweight ($1650$ lines of code), modular, and interoperable with existing repository-level translation systems.
We evaluate the effectiveness of MatchFixAgent for repository-level translation validation and repair against four existing techniques (Ibrahimzada et al., 2025; Zhang et al., 2025; Wang et al., 2025a; Ou et al., 2024). Our benchmark comprises 2,2192{,}219 source–translation function pairs, which cover $6$ PL pairs, drawn from $24$ real-world projects totaling over $900K$ lines of code (§4.1). For each translation pair, we obtain an equivalence verdict (validation outcome) from both MatchFixAgent and other techniques. Overall, MatchFixAgent returns a verdict for $99.2\%$ of pairs, while alternative approaches do so for only $71.6\%$4.2.1). On the 1,5711{,}571 pairs where both produce verdicts, MatchFixAgent agrees with other approaches in $72.8\%$ of cases. For the remaining disagreements, a systematic manual investigation finds MatchFixAgent to be correct in $60.7\%$ of cases and incorrect in the rest (§4.2.2). In translation repair, MatchFixAgent can fix $50.6\%$ of translation bugs, $32.1\%$ more than existing approaches (§4.3). We show that MatchFixAgent is compatible with different LLMs and agent frameworks, producing comparable results (§4.4). Lastly, our ablation study shows that removing code analyses and in-the-loop test generation reduces verdict accuracy by $42.3\%$, while increasing token usage by $5.2\%$4.5). These results confirm that MatchFixAgent is a viable alternative to prior work’s validation and repair approaches, while being vastly easier to adapt to new PL pairs.
Our notable contributions are:
  1. (1)
    We present MatchFixAgent, a PL-agnostic, agentic approach for validation and repair of repository-level code translation.
  2. (2)
    We demonstrate that MatchFixAgent is a viable alternative to prior work’s validation approach, while being vastly easier to adapt to new PL pairs.
  3. (3)
    We demonstrate the benefit of MatchFixAgent’s multi-agent architecture compared to simpler standalone-agent design.
2. Limitations of Prior Work
Figure 1. Illustration of key limitations of existing techniques in validation and repair of repository-level code translation and MatchFixAgent addressing them.
To demonstrate the limitations of existing techniques (Ibrahimzada et al., 2025; Zhang et al., 2025; Wang et al., 2025a; Shetty et al., 2024) for validation and repair, we use the examples in Figure 1.
Limitation 1: Generalization. Existing automated translation tools require substantial engineering effort to build validation systems for individual language pairs. For instance, Oxidizer (Zhang et al., 2025) and Syzygy (Shetty et al., 2024) require dynamic analysis and I/O extraction from source code ( a1). Oxidizer further requires a predefined feature map ( a2), which together with their dynamic analyzer is 19,00019{,}000 lines of code. Other tools like AlphaTrans (Ibrahimzada et al., 2025) and Skel (Wang et al., 2025a) validate programs using high quality type map ( a3), which requires manual maintenance over time as PLs evolve. In comparison, MatchFixAgent uses only language-agnostic LLM prompts, an off-the-shelf LLM coding agent tool, and lightweight static analysis. While the static analysis must be implemented for each PL, each PL requires approximately $280$ additional lines of code to support, hence it is extremely easy to generalize to many PLs.
Limitation 2: Validation. Most prior work depends on existing source project tests for translation validation, but these tests may have insufficient coverage. The test suites in AlphaTrans (Ibrahimzada et al., 2025) have an average of $56.57\%$ method coverage. Even when developer-written tests provide adequate coverage, they may miss the edge cases that expose subtle semantic differences between source and target languages. Consider the real-world case from Oxidizer (Zhang et al., 2025), where a Go program ( b1) that counts characters in a string is translated to Rust ( b2) using the .len() method, which counts bytes rather than characters. Oxidizer validates this pair as functionally equivalent because it only exercises this program with ASCII inputs. MatchFixAgent, in contrast, marks the translation as not equivalent, and synthesizes a test with Unicode inputs (e.g., U+1F60A, a four-byte emoji representing a single character) to confirm the inequivalence. Upon detecting the translation bug, it automatically generates a patch that uses .chars().count() to ensure proper handling of both ASCII and Unicode characters ( b3).
Figure 2. Overview of MatchFixAgent.
Limitation 3: Repair. Current translation repair techniques solely rely on simple feedback-driven approaches, which have proven inadequate in practical settings. Skel (Wang et al., 2025a), Oxidizer (Zhang et al., 2025), and Syzygy (Shetty et al., 2024) utilize multi-turn iterative prompting techniques. AlphaTrans (Ibrahimzada et al., 2025) adopts an execution trace-based reprompting strategy. For instance, consider scenario ( c1), where running the test method sequentially invokes methods m1, m2, and m3. If a translation bug is present in fragment m2, AlphaTrans reprompts all executed fragments individually without considering inter-fragment dependencies. This approach, while potentially resolving the bug in m2, risks introducing new functional bugs in previously correct fragments, such as m1. To overcome this limitation, MatchFixAgent uses code analysis and an LLM agent  ( c2). The analyses expose dependencies such as the one between m1 and m2 and the agent interacts with the execution environment to execute, validate, and iteratively refine its generated patches.
3. MatchFixAgent
In this section, we discuss MatchFixAgent and its three primary components in more detail. We first provide a high-level overview, then go into more details.
3.1. Overview
1
2
3
4
5
Input : sourceProject, sourceFunc, translatedProject, translatedFunc, LLM, tools, timeout
Output : validationRepairReport
6
7transPair $\leftarrow$ [ sourceFunc, translatedFunc ]
8
9Function async semAnalyzer (transPair, LLM):
10  cfgSrc, dfSrc $\leftarrow$ build_cfg (sourceFunc)
11  cfgTgt, dfTgt $\leftarrow$ build_cfg (translatedFunc)
12  return
13     { controlFlowAnalyzer (cfgSrc, cfgTgt, transPair, LLM),
14     dataFlowPathAnalyzer (dfSrc, dfTgt, transPair, LLM),
15     ioAnalyzer (transPair, LLM),
16     libraryAnalyzer (transPair, LLM),
17     exceptionAnalyzer (transPair, LLM),
18     specAnalyzer (transPair, LLM) }
19
20
21await semAnalysis $\leftarrow$ semAnalyzer (cfgSrc, dfSrc, cfgTgt, dfTgt)
22 testRepair $\leftarrow$ testGenRepairAgent (prompt, LLM, tools, timeout)
23 verdict $\leftarrow$ verdictAgent (semAnalysis, testRepair, LLM)
24 validationRepairReport $\leftarrow$ semAnalysis $\cup$ testRepair $\cup$ verdict
return validationRepairReport
Algorithm 1 MatchFixAgent
Figure 2 gives an overview of MatchFixAgent, which consists of three main components: (1) the Semantic Analyzer (§3.2), (2) the Test Generator & Repair Agent (§3.3), and (3) the Verdict Agent (§3.4). MatchFixAgent’s main inputs are: a translation pair (a source function and its translation), the source project, and the source project’s translation. MatchFixAgent’s outputs are: an equivalence verdict, a natural language report, and an optional patch that repairs the translation if the translation was found to be not equivalent. In addition, MatchFixAgent is configured with an LLM, a set of tools for the agents, and a timeout.
Details of MatchFixAgent’s algorithm are shown in Algorithm 1. First, the Semantic Analyzer is executed, which uses an LLM to analyze different semantic properties of the source function and its translation. We decompose this task into six sub-tasks for the LLM, which are executed independently. Each task analyzes a different semantic property of the source and translation, namely control flow (§3.2.1), data flow (§3.2.2), input and output mapping (§3.2.3), library API usage (§3.2.4), exception and error handling (§3.2.5), and specifications (§3.2.6). This sub-task architectures help to keep the LLM focused, and improves overall reliability and consistency (Anthropic, 2025). Each sub-task outputs a report that summarizes differences between the source and translation for the given semantic property.
These reports are then fed to the Test Generator and Repair Agent, which uses an off-the-shelf LLM coding agent, such as Claude Code (Team, 2025c) or Codex (Team, 2025k), to write executable test cases that may reveal inequivalent behavior. If the agent discovers inequivalent behavior, the agent also attempts to write a patch to repair the translation. This component outputs a boolean equivalence verdict, a set of tests for both the source project and translation, and an optional patch.
Finally, the Verdict Agent takes the outputs from the Semantic Analyzer and Test Generator and Repair Agent, which validates the claims made in those outputs, and provides a final verdict on the functional equivalence between the source and translation. The output includes a final boolean equivalence verdict, an overall summary, and the outputs from the Test Generator and Repair Agent.
3.2. Semantic Analyzer
The Semantic Analyzer takes as input the translation pair and an LLM. It first computes a control flow graph (CFG) and data flow graph (DFG) (described in Sections 3.2.1 and 3.2.2 respectively), and then calls six sub-analyzers in parallel, each of which analyzes a different semantic property of the translation pair. Each sub-analyzer invokes the LLM with a custom prompt describing the analysis to perform. The prompts are relatively simple and short. The prompt first defines a role for the LLM (‘‘You are an expert in...’’), a general definition of functional equivalence, and a specific definition of equivalence for the semantic property. It then instructs the analyzer to output an equivalence verdict and explanation for the specific semantic property it analyzed. In addition, certain analyzers output examples to demonstrate inequivalence. The final output of the Semantic Analyzer is a 6-tuple containing a JSON-formatted output of each sub-analyzer. The following subsections provide more details on the prompts of the six sub-analyzers.
============= Source Code ============
def max(a, b): """
The function max(a, b) returns the greater of the two input values a and b. If a is greater than b, it returns a; otherwise, it returns a.
"""    if a > b:
    return a
   else:
    return b
========= Control Flow Graph =========
Node 0: PARAMETERS
   parameter: a, parameter: b
   $\rightarrow$ Node 1: ENTRY
Node 1: ENTRY
   $\rightarrow$ Node 2: if a > b:
Node 2:
   if a > b:
   $\rightarrow$ Node 3 [label="a > b"]:
     return a
   $\rightarrow$ Node 4 [label="(a <= b)"]:
     return b
=========== Data Flow Paths ===========
Variable: a
   Path 1:
    CFG Node Path: 0 $\rightarrow$ 1 $\rightarrow$ 2
   Path 2:
    CFG Node Path: 0 $\rightarrow$ 1 $\rightarrow$ 2 $\rightarrow$ 3
Variable: b
Path 1:
    CFG Node Path: 0 $\rightarrow$ 1 $\rightarrow$ 2
Path 2:
    CFG Node Path: 0 $\rightarrow$ 1 $\rightarrow$ 2 $\rightarrow$ 4
Figure 3. CFG and DFP structures extracted by the Semantic Analyzer component in MatchFixAgent.
3.2.1. Control Flow Analyzer
The Control Flow Analyzer is prompted to analyze the the control flow structures of the source and translation are equivalent, looking for inequivalences like reordered conditions, missing branches, or altered loop termination criteria. To aid this task, we provide the LLM with textual representations of the source and translation’s CFGs. An example is shown in Figure 3. To compute the CFG, we use Tree-Sitter (Tree-Sitter, 2025) to construct an abstract syntax tree of the function, and then extract basic blocks and control flow structures. Tree-sitter supports 165165+ PLs, making this process PL-agnostic. Each PL supported by MatchFixAgent required approximately $280$ lines of code, making it very easy to support many PLs.
1
2
3
4
5
Input : cfgSource, cfgTarget, fragments, model
Output : cfgAnalysis
6
7Function abstractGraph (cfg):
8  for ($u$, $v$) with edgeedge $\in$ cfg do
9     uType, vType $\leftarrow$ classifyNode ($u$), classifyNode ($v$)
10     edgeType $\leftarrow$ classifyEdge (edgeedge)
11     nodes $\leftarrow$ nodes $\cup$ uType $\cup$ vType
12     edges $\leftarrow$ edges $\cup$ $\langle$ uType, edgeType, vType $\rangle$
13 return nodes, edges
14
15sourceNodes, sourceEdges $\leftarrow$ abstractGraph (cfgSource)
16targetNodes, targetEdges $\leftarrow$ abstractGraph (cfgTarget)
17nodeSim $\leftarrow$ jaccardSimilarity (sourceNodes, targetNodes)
18edgeSim $\leftarrow$ jaccardSimilarity (sourceEdges, targetEdges)
19similarity $\leftarrow$ ( $0.5$ $\times$ nodeSim) ++ ( $0.5$ $\times$ edgeSim)
20
21if similarity $\geqslant$ threshold == $0.7$ then
22  cfgAnalysis $\leftarrow$ { "is_equivalent": "yes" }
23else
24  cfgAnalysis $\leftarrow$ LLM (cfgSource, cfgTarget, fragments, model)
return cfgAnalysis
Algorithm 2 Control Flow Analyzer
To improve reliability and reduce costs, the control flow analyzer first (symbolically) computes a similarity score between the CFGs, and, if it falls above a threshold, it immediately returns an equivalent verdict without invoking the LLM. The overall procedure is shown in Algorithm 2. The analyzer abstracts each graph into canonical forms (lines 1199) capturing node types (e.g., conditionals, loops, exception handlers) and edge types (control transfer relationships), then computes the structural similarity score based on the Jaccard index (Cha, [n. d.]) (lines 10101212). We use $0.7$ as the threshold. This results in approximately $25\%$ of LLM invocations being skipped in our experiments, making this threshold relatively stringent.
3.2.2. Data Flow Analyzer
The Data Flow Analyzer is prompted to evaluate whether the flow of data within the source and translation is equivalent, looking for issues like unused variables. To aid the LLM, we provide textual representations of the source and translation’s data flow graphs (DFGs). An example is shown in Figure 3. We keep our data flow computation extremely simple. For each statement in the AST, we extract variable names, label them as a def or a use, and associate them with a CFG node. We then extract def-use chains. This is primarily a syntactic analysis. We do not handle challenging problems such as aliasing, concurrency, or context sensitivity. The per-PL implementation effort is approximately $280$ lines of code based on the six PLs supported by MatchFixAgent.
Similar to our control flow analyzer, we compute a similarity score between the DFGs, and short-circuit if it falls above a threshold. The analyzer, show in Algorithm 3, first extracts def-use chains for parameters and local variables, capturing how data values are defined, propagated, and consumed. The extracted paths are compared using edit distance (Miller et al., 2009) as the similarity measure (lines 111313). We again use $0.7$ as the threshold, which results in approximately $35\%$ of LLM invocations being skipped.
1
2
3
4
5
Input : dfSource, dfTarget, fragments, model
Output : dfAnalysis
6
7
8
9Function computeEditDistance (srcPath, tgtPath):
10  sim_a $\leftarrow 0$
11  foreach xPathxPath $\in$ srcPath do
12     best $\leftarrow$ 0
13     foreach yPathyPath $\in$ tgtPath do
14        score $\leftarrow$ jaccardSimilarity (xPathxPath, yPathyPath)
15        best $\leftarrow$ max (best, score)
16    sim_a $\leftarrow$ sim_a ++ best
17 sim_a $\leftarrow$ sim_a // $|$ srcPath $|$
18  sim_b $\leftarrow$ "repeat loop with srcPath and tgtPath swapped"
19  return (sim_a ++ sim_b) // 22
20
21srcPath, tgtPath $\leftarrow$ getVariablePaths (dfSource, dfTarget)
22similarity $\leftarrow$ computeEditDistance (srcPath, tgtPath)
23if similarity $\geqslant$ threshold == $0.7$ then
24  dfAnalysis $\leftarrow$ { "is_equivalent": "yes" }
25else
26  dfAnalysis $\leftarrow$ LLM (dfSource, dfTarget, fragments, model)
27
return dfAnalysis
Algorithm 3 Data Flow Path Analyzer
3.2.3. IO Analysis
The IO Analyzer is prompted to evaluate whether the observable input-output behavior of the source and target fragments is semantically equivalent. The prompt includes an IO equivalence definition, which assess five dimensions: (1) semantic equivalence of accepted inputs, (2) consistency of produced outputs, (3) preservation of side effects (e.g., file operations, network calls, or global state modifications), (4) uniform handling of edge cases, and (5) similarity in performance-critical complexity. The LLM is prompted also prompted to produce a plausible input that would trigger dissimilar IO behavior if it believes the translation is inequivalent. This methodology catches inequivalences such as differing error messages, inconsistent encoding assumptions, or missing side effects—often overlooked by structural analyses (§3.2.13.2.2) alone.
3.2.4. Library Analyzer
The Library API Analyzer is prompted to consider the behavior of external library APIs called in the source and translation, and evaluate whether their differences result in inequivalent behavior. This analyzer primarily detects subtle differences between similar library APIs in the source and translation. It provides suggestions to fix inequivalent behavior as well.
3.2.5. Exception & Error Analyzer
The Exception and Error Handling Analyzer is prompted to validate whether error detection, exception raising, and error recovery mechanisms in the source and target code fragments are functionally equivalent. The prompt include five dimensions for equivalence: (1) detecting and handling the same error conditions, (2) using semantically equivalent exception/error types, (3) producing equivalent error messages or codes, (4) preserving consistent recovery mechanisms, and (5) propagating errors in equivalent ways. If neither fragment implements explicit error handling, the analyzer deems them equivalent for this dimension. Otherwise, it statically identifies exception constructs (e.g., try-catch, throw, return-error patterns) and uses LLM reasoning to compare semantics. For instance, if the source raises a specific FileNotFoundError while the target raises a generic IOException, the discrepancy is flagged, as it may affect upstream handling. Where differences exist, the analyzer also recommends target-language error handling constructs that align with the source’s semantics.
3.2.6. Specifications Analyzer
The Specification Analyzer is prompted assesses whether the source and target code fragments adhere to the same explicit or implicit functional specifications. The prompt includes Specification equivalence definition which state that the source and translation should: (1) fulfill the same documented or inferred functional requirements, (2) satisfy identical pre-conditions and post-conditions, (3) maintain the same invariants, and (4) handle the same input domain, including edge cases. The LLM is instructed to extract available specifications from function signatures, type annotations, docstrings, and relevant comments, or, when no formal documentation exists, infer behavioral contracts from code semantics. The LLM is asked to compare the contracts of the source and translation. For example, if the source specifies ‘‘returns 1 on success, 0 on failure’’ and the target returns Boolean values, the inconsistency is flagged. In such cases, the LLM is instructed to produce a formalized specification that reconciles both implementations and provides counterexamples demonstrating behavior mismatch.
3.3. Test Generator & Repair Agent
The Test Generator & Repair Agent uses an off-the-shelf LLM coding agent and the reports from the Semantic Analyzer to write and execute tests that demonstrate functional (in)equivalence. This agent helps catch hallucinations and confirm the claims in the Semantic Analyzer reports. The prompt for the agent is shown in Figure 4, which includes a definition of equivalence, instructions to write tests in both the source and target language that test the (in)equivalence of the translation, and finally instructions to repair the translation if it is not equivalent. We use Claude Code (Team, 2025c) as the agent for most of our experiments, which comes with a set of tools out of the box, namely, reading + writing files, executing arbitrary shell commands, and searching the web. The agent outputs an overall equivalence verdict, a set of tests in both the source and target language, and a translation patch if the agent believed the translation was not equivalent.
<fragment_details>
<source_fragment_details> <path to source file> <implementation of source fragment> </source_fragment_details>
<target_fragment_details> <path to target file> <implementation of target fragment> </target_fragment_details>
</fragment_details>
<instruction> You are an expert agent specializing in test generation and code repair. Based on the analysis from multiple expert agents regarding functional equivalence between \$SOURCE_LANGUAGE and \$TARGET_LANGUAGE implementations of the given method/function, your task is to generate tests and repair the target implementation, if necessary.    <functional_equivalence_definition> Two code fragments in different programming languages are considered functionally equivalent if, when executed on the same input, they always have identical program states at all corresponding points reachable by program execution, and they both produce the same output upon termination.    </functional_equivalence_definition>
<rules_and_general_notes> 1. Consider the Semantic Analyzer results ..., 2. Generate tests ... </rules_and_general_notes>
</instruction>
<semantic_analysis_results> {"control_flow": <>, "data_flow": <>, "io": <>, "lib_api": <>, ...} </semantic_analysis_results>
<final_response_format> {"is_equivalent": <>, "explanation": <>, "tests": <>, "patch": <>, ...} </final_response_format>
Figure 4. Prompt structure of Test Generator and Repair Agent.
3.4. Verdict Agent
The final component of MatchFixAgent is the Verdict Agent, which produces a definitive assessment of the translation’s correctness by synthesizing the information from the previous two stages. The Verdict Agent takes as input the semantic analysis report and the test execution + repair report. It leverages another LLM agent to consolidate these results into a final verdict. This agent’s primary job is to (1) confirm the results of the Test Generator & Repair Agent, and (2) to produce a condensed summary of the results, which is useful for end-users.
4. Evaluation
We evaluate MatchFixAgent and answer the following research questions:
  1. RQ1:
    Effectiveness of MatchFixAgent in Translation Validation. To what extent can MatchFixAgent automatically validate repository-level code translation? How does MatchFixAgent compare against existing validation systems?
  2. RQ2:
    Effectiveness of MatchFixAgent in Translation Repair. How effective is MatchFixAgent in repairing translation bugs from real-world projects? How does the repair component compare with existing tools?
  3. RQ3:
    Development Cost and Adaptability. How does the development cost and adaptability of MatchFixAgent compare to existing work? Does it work with other LLMs and agents?
  4. RQ4:
    Ablation Study. How do the semantic analyzer and test generator components contribute to MatchFixAgent’s effectiveness? Can a standalone code agent perform similarly to MatchFixAgent?
4.1. Experimental Setup
Table 1. Details of benchmarks [^3] Some projects in Skel (e.g., colorsys) are part of a bigger project (Team, 2025m). The reported Star and Fork numbers belong to that bigger project. from existing techniques used in MatchFixAgent. LoC: Lines of code in the source project.
Tool Project
Source
Language
Target
Language
Total #
Trans. Pairs
LoC
Test
Coverage (%)
Stars Forks
Oxidizer (Zhang et al., 2025) checkdigit (Tonomori, 2025) Go Rust 29 428 86.2 111 8
go-edlib (Bollon, 2025) 24 639 100 517 27
histogram (Cortex, 2025) 19 314 68.4 176 31
nameparts (Polera, 2025) 15 413 100 43 5
stats (Flynn, 2025) 53 1241 98.1 2989 170
textrank (Belicza, 2025) 52 1132 100 217 22
AlphaTrans (Ibrahimzada et al., 2025) cli (Foundation, 2025b) Java Python 273 37841 100 372 201
csv (Foundation, 2025c) 235 33072 100 392 278
fileupload (Foundation, 2025d) 192 3567 100 246 185
validator (Foundation, 2025e) 646 41605 95.5 216 164
Skel (Wang et al., 2025a) bst (Algorithms, 2025a) Python JavaScript 19 123 100 203000 47000
colorsys (Team, 2025l) 8 120 96.3 67900 32300
heapq (Team, 2025n) 22 189 78.1 67900 32300
html (Team, 2025o) 44 684 65.3 67900 32300
mathgen (Weiler, 2025) 81 735 93.8 711 183
rbt (Algorithms, 2025b) 27 366 88.1 203000 47000
strsim (Luo, 2025) 64 654 20.8 1014 125
toml (Pearson, 2025) 72 1206 62.5 1126 192
RustRepoTrans (Ou et al., 2024) charset (Jawah, 2025) Python Rust 33 4231 100 672 56
deltachat (Chat, 2025) C 125 23116 98.4 306 28
iceberg-java (Foundation, 2025f) Java 25 592793 100 7700 2700
iceberg-py (Foundation, 2025g) Python 44 49746 97.7 805 327
crypto-c (Foundation, 2025a) C 20 5922 100 36 15
crypto-java (Foundation, 2025h) Java 97 110261 100 2 7
Total 2219 910398 89.6 627351 195624
4.1.1. Benchmark.
We evaluate MatchFixAgent on benchmarks used in prior work on automated repository-level translation. Each benchmark problem is a translation pair: the source function and the corresponding translation. The task for each benchmark problem is to give a verdict on the functional equivalence of pairs in two different PLs, and repair translation in case of equivalence. Our subjects are open-source repository-level translations with equivalence verdicts available [^4] We exclude rule-based transpilers as they do not propose a validation mechanism, i.e., their translation is (theoretically) correct by construction. from the peer-reviewed literature [^5] This criterion excludes techniques such as RustMap (Cai et al., 2025) and C2SaferRust (Nitin et al., 2025).. We make selections across a diverse set of PL pairs.
Table 1 summarizes our subject translation pairs. We collect subjects from three recent repository-level code translation techniques (Ibrahimzada et al., 2025; Zhang et al., 2025; Wang et al., 2025a). We did not include Syzygy (Shetty et al., 2024) because its artifact does not provide a validation system for individual functions (it only provides end-to-end tests). These works generated translations of real-world open source GitHub projects, and performed equivalence validation at the individual function level. Since AlphaTrans (Ibrahimzada et al., 2025) is evaluated on a large number of functions, we randomly sample $1346$ of its total $4643$ translation pairs [^6] AlphaTrans translated both application and test code. In this work, we only included the application code translation pairs. While reviewing their artifacts, we also noted $11$ translation pairs to be dead code and excluded them..
To demonstrate the adaptability of MatchFixAgent to more PL pairs, we also collect subjects from RustRepoTrans (Ou et al., 2024), a benchmark consisting of human-written translations into Rust and unit tests. We exclude translation pairs collected from the libp2p (libp2p, 2025) projects in RustRepoTrans due to the presence of non-deterministic flaky tests that may result in false negatives, i.e., functional inequivalence while translation is correct, unfairly biasing comparison in favor of MatchFixAgent. In total, we collect 2,2192{,}219 translation pairs with over $900K$ lines of code from $24$ projects and in $6$ different PL pairs.
4.1.2. LLMs.
Major software engineering leaderboards (bench Team, 2025; Team, 2025b) have shown that Claude Sonnet (Team, 2025a) outperforms other proprietary LLMs, such as OpenAI GPT-4o (Team, 2025j) and Google Gemini Pro (Team, 2025e). Therefore, we use Anthropic’s Claude 3.7 Sonnet (Team, 2025a) and Claude Code (1.0.511.0.51(Team, 2025c) as the main LLM and agent in all our experiments. To show the adaptability of MatchFixAgent to different LLMs and agentic frameworks, we repeat a subset of experiments using OpenAI o4-mini-2025-04-16 (OpenAI, 2025) and Codex (Team, 2025k)4.4). To support future research and external validation, MatchFixAgent logs the inputs, intermediate agent interactions, tool execution results, and outputs of the LLM, and supports visualizing and inspecting these logs. MatchFixAgent terminates within the budget of 1,0001{,}000 seconds. We empirically set this timeout after analyzing the execution time of $300$ samples.
4.1.3. Competing Validation & Repair Tools
We compare MatchFixAgent’s validation technique with the automated validation techniques proposed by Skel (Wang et al., 2025a), Oxidizer (Zhang et al., 2025), and AlphaTrans (Ibrahimzada et al., 2025). Except RustRepoTrans, other approaches do not explicitly report the repair results, as the repair process is interleaved with translation in the loop. For those techniques (Ibrahimzada et al., 2025; Zhang et al., 2025; Wang et al., 2025a), we compare MatchFixAgent repair results with their final translation success.
4.1.4. MatchFixAgent Implementation.
We implement our structure-based semantic analysis on top of Tree-Sitter (Tree-Sitter, 2025), as it supports $165+$ languages, including six PLs we target in this study. For running tests and validating patches, MatchFixAgent uses Rust $1.87.0$ (Team, 2025q), Python $3.12.9$ (Team, 2025p), Java $21.0.7$ (Team, 2025g), Node $22.16.0$ (Team, 2025i), GCC $7.3.1$ (Team, 2025d), and Go $1.24.4$ (Team, 2025f).
4.2. RQ1: Effectiveness of MatchFixAgent in Translation Validation
Table 2. Effectiveness of MatchFixAgent in translation validation compared to existing techniques. EQ: Equivalent, NEQ: Not Equivalent, VF: Validation Failure, Agreement: number and percentage of translation pairs where MatchFixAgent’s and the existing tool’s verdicts agree, Disagreement: percentage of disagreements ruled in favor of Tool and MatchFixAgent (Ours). VFs are excluded from Agreement and Disagreement calculations.
Tool Project Total # Trans. Pairs Tool Validation MatchFixAgent Agreement Disagreement
EQ NEQ VF EQ NEQ VF Tool Ours
Oxidizer checkdigit 29 21 (72.4) 8 (27.6) 0 (0) 22 (75.9) 7 (24.1) 0 (0) 24 (82.8) 0.0 100
go-edlib 24 18 (75) 6 (25) 0 (0) 16 (66.7) 8 (33.3) 0 (0) 14 (58.3) 11.1 88.9
histogram 19 12 (63.2) 7 (36.8) 0 (0) 11 (57.9) 7 (36.8) 1 (5.3) 8 (44.4) 20.0 80.0
nameparts 15 9 (60) 6 (40) 0 (0) 12 (80) 3 (20) 0 (0) 12 (80) 33.3 66.7
stats 53 38 (71.7) 14 (26.4) 1 (1.9) 37 (69.8) 16 (30.2) 0 (0) 35 (67.3) 0.0 100
textrank 52 40 (76.9) 12 (23.1) 0 (0) 34 (65.4) 18 (34.6) 0 (0) 28 (53.8) 42.9 57.1
[…]
Alpha Trans cli 273 210 (76.9) 24 (8.8) 39 (14.3) 210 (76.9) 60 (22) 3 (1.1) 176 (76.2) 30.0 70.0
csv 235 97 (41.3) 61 (26) 77 (32.8) 185 (78.7) 49 (20.9) 1 (0.4) 108 (68.8) 30.0 70.0
fileupload 192 19 (9.9) 1 (0.5) 172 (89.6) 144 (75) 48 (25) 0 (0) 16 (80) 25.0 75.0
validator 646 247 (38.2) 103 (15.9) 296 (45.8) 483 (74.8) 163 (25.2) 0 (0) 225 (64.3) 20.0 80.0
[…]
Skel bst 19 19 (100) 0 (0) 0 (0) 14 (73.7) 5 (26.3) 0 (0) 14 (73.7) 20.0 80.0
colorsys 8 8 (100) 0 (0) 0 (0) 7 (87.5) 1 (12.5) 0 (0) 7 (87.5) 0.0 100
heapq 22 19 (86.4) 3 (13.6) 0 (0) 12 (54.5) 10 (45.5) 0 (0) 13 (59.1) 50.0 50.0
html 44 40 (90.9) 2 (4.5) 2 (4.5) 35 (79.5) 9 (20.5) 0 (0) 33 (78.6) 66.7 33.3
mathgen 81 77 (95.1) 4 (4.9) 0 (0) 65 (80.2) 16 (19.8) 0 (0) 67 (82.7) 50.0 50.0
rbt 27 26 (96.3) 0 (0) 1 (3.7) 23 (85.2) 4 (14.8) 0 (0) 22 (84.6) 75.0 25.0
strsim 64 50 (78.1) 0 (0) 14 (21.9) 56 (87.5) 8 (12.5) 0 (0) 44 (88) 40.0 60.0
toml 72 37 (51.4) 10 (13.9) 25 (34.7) 49 (68.1) 22 (30.6) 1 (1.4) 33 (71.7) 40.0 60.0
[…]
RustRepo Trans charset 33 20 (60.6) 13 (39.4) 0 (0) 14 (42.4) 19 (57.6) 0 (0) 25 (75.8) 75.0 25.0
deltachat 125 54 (43.2) 69 (55.2) 2 (1.6) 39 (31.2) 84 (67.2) 2 (1.6) 92 (76) 90.0 10.0
iceberg-java 25 9 (36) 16 (64) 0 (0) 3 (12) 16 (64) 6 (24) 15 (78.9) 100 0.0
iceberg-py 44 15 (34.1) 28 (63.6) 1 (2.3) 11 (25) 29 (65.9) 4 (9.1) 35 (89.7) 100 0.0
crypto-c 20 16 (80) 4 (20) 0 (0) 7 (35) 13 (65) 0 (0) 11 (55) 66.7 33.3
crypto-java 97 39 (40.2) 58 (59.8) 0 (0) 30 (30.9) 67 (69.1) 0 (0) 86 (88.7) 100 0.0
[…]
To answer this RQ, we run MatchFixAgent on each translation pair to obtain an equivalence verdict, and compare it with the verdict of existing tools (§4.2.1). We then investigate disagreements between verdicts to determine which verdict is correct (§4.2.2).
4.2.1. Translation Validation
The columns under Tool Validation and MatchFixAgent in Table 4.2 summarize the equivalence verdicts for the competing validation tool and MatchFixAgent, respectively. There are three types of equivalence verdicts: (1) Equivalent (EQ) indicating the source function and its translation are equivalent, (2) Not Equivalent (NEQ) indicating they are inequivalent, and (3) Validation Failure (VF) indicating that the tool failed to provide a verdict. Competing tools can fail to provide a verdict if (1) the source project did not have unit tests covering the function, or (2) the competing tool’s language interoperability mechanism crashes before providing verdict. MatchFixAgent can fail to provide a verdict if the timeout limit is reached. The Agreement column shows the number and percentage where both MatchFixAgent’s and others verdicts agree or disagree. The Disagreement column shows the result of our human investigation on disagreements. The Tool sub-column shows the percentage of disagreements where the existing tool’s verdict was correct, and Ours sub-column shows the same metric for MatchFixAgent. On average, MatchFixAgent takes $309$ seconds to produce a verdict, with an average cost of \$$1.22$ and a total cost of \$2,710.452{,}710.45.
These results demonstrate effectiveness of MatchFixAgent in providing an equivalence verdict: MatchFixAgent provides verdicts for $99.7\%$ of translation pairs from AlphaTrans and Skel, whereas these techniques report verdicts for only $56.6\%$ and $87.5\%$ of their studied translation pairs, respectively. In addition, MatchFixAgent’s verdicts show a high level of agreement with all prior work, ranging from $63.7\%$ to $80.2\%$.
4.2.2. Analysis of the Disagreements
Both MatchFixAgent and existing validation approaches are prone to false positives (i.e. the verdict is equivalent but the translation is not) and false negatives. To determine false positives and false negatives, we perform a manual investigation.
We first categorize disagreements into two cases: $D_{1}$, where MatchFixAgent produced a not equivalent verdict and the other disagreed; and $D_{2}$, where MatchFixAgent produced an equivalent verdict and the other disagreed. Due to large number of studied translation pairs, we randomly sample five instances of both $D_{1}$ and $D_{2}$ from each of the $24$ projects. If a project had fewer than five disagreements, we considered all instances without sampling. Two authors [^7] The selected authors have total experience of $17$ years in academia and $4.5$ years in industry. independently reviewed disagreements to verify whether the not equivalent verdict was correct (by MatchFixAgent in $D_{1}$ and by others in $D_{2}$). If the not equivalent verdict is correct, the reviewer rules it in favor of the tool that said not equivalent, otherwise they rule it favor of the tool that said equivalent. During the investigation, 3 disagreements with Oxidizer were due to the tool’s function mocking not being enabled, and were unable to enable it. In addition, 11 disagreements with RustRepoTrans were due to the correct translation not being ‘‘1:1’’, or in other words, the correct translation is not functionally equivalent to the source function. These disagreements would have been unfairly ruled in favor of MatchFixAgent, and so were filtered out, leaving us with $145$ disagreement cases ($D_{1}=92$, $D_{2}=53$). When the reviewer’s resolution conflicted ($18.6\%$ of cases), they met with each other and agreed on the final resolution.
The Disagreement columns in Table 4.2 summarize the result of our human investigation. The Tool column shows the percentage of disagreements ruled in favor of the existing validation tool, and the Ours columns shows the percentage ruled in favor of MatchFixAgent. The disagreement resolutions show that MatchFixAgent ’s verdicts are often more accurate than existing automated validation tools. MatchFixAgent’s verdicts are significantly more accurate than Oxidizer and AlphaTrans— disagreements are ruled in favor of MatchFixAgent in $84.1\%$ and $73.5\%$ of cases, respectively. Compared to Skel, MatchFixAgent shows similar accuracy ($53.5\%$). On RustRepoTrans, MatchFixAgent’s accuracy fairs worse ($12.5\%$), which is due to the relative complexity of its translation pairs.
Existing validation approaches produced $88$ incorrect verdicts, $80$ of which fell into three categories: (1) Inadequate Unit Tests (42 of cases), (2) Excessively Strict Equivalence Definition (11 of cases), and (3) Language Interoperability Bug (27 of cases). A language interoperability bug means that the tool’s process for converting concrete inputs in the source language to the target language did not preserve equivalence. For example, Oxidizer’s conversion of Go runes (which represent a Unicode character) to Rust chars resulted in different Unicode characters.
On the benchmarks associated with automated validation tools (Skel, Oxidizer, AlphaTrans), MatchFixAgent produced 36 incorrect verdicts, $34$ of which fell into three categories: (1) Hallucination (23 cases), (2) Inadequate Unit Tests (4 cases) (meaning MatchFixAgent missed an input that would demonstrate inequivalence), (3) Infeasible Input (7 cases). An infeasible input means that the LLM discovered an input where the source function and translation produce different outputs, but the input can never occur when the project is used as intended. Infeasible inputs often involve directly initializing private class/struct members, or calling private helper methods in unintended ways.
On RustRepoTrans’s benchmarks, MatchFixAgent produced $21$ incorrect verdicts. Two of the main causes are similar to the other benchmarks: (1) Hallucination (7 cases) and (2) Inadequate Unit Tests (6 cases). The increased rates of these two causes are due to the relative complexity and size of RustRepoTrans’s projects. The other major cause is Excessively Strict Equivalence Definition (6 cases). As previously mentioned, RustRepoTrans’s translations include refactors to make the translation more idiomatic, which creates ambiguity around the proper definition equivalence.
4.2.3. Representative Examples of Incorrect Verdicts
The following code snippet shows an example of Inadequate Unit Tests on a translation pair from Oxidizer. The functions both calculate the edit distance between two strings. The functions are not equivalent because Go’s len() function counts Unicode characters, whereas Rust’s .len() function counts bytes. Oxidizer incorrectly validates the Rust translation as equivalent because the source project’s unit tests do not cover non-ASCII inputs. However, MatchFixAgent successfully generates a test with non-ASCII inputs demonstrating they are not equivalent.
——————- GO SOURCE CODE ——————-
func LCSEditDistance(str1 string, str2 string) int {
[…]
lcs := LCS(s1, s2)
return (len([]rune(s1)) - lcs) + (len([]rune(s2)) - lcs)
—————— RUST TRANSLATION ——————
pub fn lcs_edit_distance(str1: &str, str2: &str) -> Result<i32> {
// if conditions
let lcs_len = lcs(str1, str2)?;
let edit_distance = (str1.len() as i32 - lcs_len) + (str2.len() as i32 - lcs_len);
Ok(edit_distance)
The next code snippet demonstrates an example of an Infeasible Input taken from the textrank project. The below Go function inserts a value into the map ranks.SentenceMap (respectively, ranks.sentence_map for the Rust translation). MatchFixAgent discovers that these functions return different values when the map is initially { 1 : ‘‘SomeString’’ } (the Go returns 0 while the Rust returns 1). However, when the textrank is used properly via its public interface, this initial state for the map cannot occur. The map will always contain keys from $1$ to $n$, where $n$ is the number of map entries. Under this precondition, the functions are equivalent.
——————- GO SOURCE CODE ——————-
func addSentence(ranks *Rank, sentence ParsedSentence) int {,
ranks.SentenceMap[len(ranks.SentenceMap)] = sentence.GetOriginal()
return len(ranks.SentenceMap) - 1
[…]
—————— RUST TRANSLATION ——————
pub(crate) fn add_sentence(ranks: &mut Rank, sentence: ParsedSentence) -> Result<i32, Error> {
let sentence_id = ranks.sentence_map.len() as i32;
ranks.sentence_map.insert(sentence_id, sentence.original.clone());
Ok(sentence_id)
The next code snippet demonstrates an example of an Excessively Strict Equivalence Definition taken from the deltachat-core project. Both the C and Rust function retrieve a field blobdir. MatchFixAgent states that these are not equivalent because (1) the Rust function does not perform a null check, and (2) the C function returns a copy of blobdir whereas the Rust returns a reference. While this is true, the translation follows Rust’s idioms, and a developer would not care about these differences. Rust’s type system prevents blobdir from ever being null, and the Rust translation returns an immutable reference, preventing the caller from modifying the return value.
——————- C SOURCE CODE ——————–
char* dc_get_blobdir(const dc_context_t* context) {
if (context==NULL $|$ context->magic!=DC_CONTEXT_MAGIC) {
return dc_strdup(NULL);
return dc_strdup(context->blobdir);
—————— RUST TRANSLATION ——————
pub fn get_blobdir(&self) -> &Path {
[…]
Summary. Compared with existing automated validation approaches (Skel, Oxidizer, and AlphaTrans), MatchFixAgent is more reliable at producing equivalence verdicts. MatchFixAgent provides verdicts for 99.299.2% of benchmarks, compared to $71.6\%$ for existing approaches. In addition, MatchFixAgent’s verdicts are more accurate than existing approaches. MatchFixAgent’s verdicts agree on $72.8\%$ of benchmarks, and on $60.7\%$ of disagreements, human investigation shows MatchFixAgent’s verdict is correct. On the more challenging benchmarks of RustRepoTrans, MatchFixAgent’s verdict still show a high agreement rate, but the disagreement cases reveal for improvement.
4.3. RQ2: Effectiveness of MatchFixAgent in Translation Repair
We investigate the effectiveness of MatchFixAgent in automated repair of translation bugs, and its ability to improve code coverage of existing projects. Table 3 shows the results of this research question. To fairly evaluate the effectiveness of existing tools and MatchFixAgent in translation repair, we extract a subset of translations where both techniques generated a patch.
In total, $\frac{265}{2219}$ ($11.9\%$) buggy translations were considered for our study. To validate patches, we used original project tests that previously failed on buggy translations, and only considered a patch correct when all failing tests passed. We did not use generated tests by MatchFixAgent to avoid bias in our evaluation. Column Tool Repaired shows the number of buggy translations repaired by existing techniques. Only $\frac{49}{265}$ ($18.5\%$) bugs have been repaired by prior techniques. Except for RustRepoTrans (Ou et al., 2024), other code translation tools failed to generate correct patches to repair translation bugs. Skel (Wang et al., 2025a) reprompts the same LLM for repairing bugs, but then requires a user to manually provide a fix. AlphaTrans (Ibrahimzada et al., 2025) and Oxidizer (Zhang et al., 2025) generates patches in the loop, however, no effectiveness were reported in their papers, and we could not repair any translation bugs using their tools. Moreover, patches by AlphaTrans and Oxidizer could not be validated mostly because of limitations in their validation system. For example, the following code snippets show instance 11 from the project commons-validator in AlphaTrans. The GraalVM-based validation system in AlphaTrans does not validate this translation as functionally equivalent, although our manual investigation indicates that the Python translation is correct. Therefore, the limitation in AlphaTrans is mostly due to its validation system being unable to validate LLM patches.
Table 3. Effectiveness of MatchFixAgent in translation repair compared against existing techniques. NEQ: Not Equivalent, NR: Not Reported/Repaired.
Tool Project
# Total
Trans. Pair
Tool NEQ $\cap$
MatchFixAgent NEQ
Tool
Repaired
MatchFixAgent
Repaired
Disagreement
Repaired
Coverage
(Improvement) %
Oxidizer checkdigit 29 5 (17.2) NR 5 (100) 2 (100) 86.2 (0)
go-edlib 24 2 (8.3) NR 1 (50) 4 (100) 100 (0)
histogram 19 2 (10.5) NR 1 (50) 2 (66.7) 68.4 (0)
nameparts 15 3 (20) NR 1 (33.3) 0 (0) 100 (0)
stats 53 6 (11.3) NR 6 (100) 5 (100) 100 ($\uparrow$1.9)
textrank 52 3 (5.8) NR 3 (100) 2 (100) 100 (0)
[…]
Alpha Trans cli 273 9 (3.3) NR 7 (77.8) 3 (100) 100 ($\uparrow$6.2)
csv 235 20 (8.5) NR 16 (80) 2 (100) 100 ($\uparrow$11.9)
fileupload 192 0 (0) - - 2 (100) 98.9 ($\uparrow$78.1)
validator 646 34 (5.3) NR 23 (67.6) 3 (100) 99.3 ($\uparrow$36.8)
[…]
Skel bst 19 0 (0) - - 4 (100) 100 (0)
colorsys 8 0 (0) - - 1 (100) 100 (0)
heapq 22 2 (9.1) NR 1 (50) 3 (100) 100 (0)
html 44 1 (2.3) NR 0 (0) 2 (100) 100 ($\uparrow$4.5)
mathgen 81 3 (3.7) NR 1 (33.3) 2 (66.7) 100 (0)
rbt 27 0 (0) - - 1 (100) 100 ($\uparrow$3.7)
strsim 64 0 (0) - - 3 (100) 100 ($\uparrow$21.9)
toml 72 5 (6.9) NR 3 (60) 3 (100) 100 ($\uparrow$34.7)
[…]
RustRepo Trans charset 33 12 (36.4) 5 (41.7) 7 (58.3) 1 (100) 100 (0)
deltachat 125 60 (48) 10 (16.7) 11 (18.3) 1 (100) 100 ($\uparrow$1.6)
iceberg-java 25 12 (48) 1 (8.3) 1 (8.3) 0 (0) 100 (0)
iceberg-py 44 25 (56.8) 4 (16) 6 (24) 0 (0) 100 ($\uparrow$2.3)
crypto-c 20 4 (20) 1 (25) 2 (50) 1 (100) 100 (0)
crypto-java 97 57 (58.8) 28 (49.1) 39 (68.4) 0 (0) 100 (0)
[…]
—————— JAVA SOURCE CODE ——————
public boolean isOn(long flag) {
return (this.flags & flag) == flag;
—————– PYTHON TRANSLATION —————–
def isOn(self, flag: int) -> bool:
[…]
Column MatchFixAgent Repaired shows the number of translation bugs successfully repaired by our approach. In total, MatchFixAgent repaired $\frac{134}{265}$ ($50.6\%$) of translation bugs, $32.1\%$ more than existing reprompting-based techniques. Given the limitations in validation system of AlphaTrans discussed earlier, we manually investigate and validate patches from this tool. The following code snippets demonstrate instance 276 from the project commons-validator in AlphaTrans which its reprompting-based repairing could not generate a correct patch. By contrast, MatchFixAgent successfully repairs this translation bug with the help of its Library Analyzer (§3.2.4). The report generated by this semantic analyzer indicate "... the standard Python datetime.date class does not have a SHORT attribute ..." which is correct. The Test Generator & Repair Agent (§3.3) then leverages this analysis and successfully generate a patch by replacing SHORT with constant 3.
—————— JAVA SOURCE CODE ——————
public static DateValidator DateValidator1() {
return new DateValidator(true, DateFormat.SHORT);
—————– PYTHON TRANSLATION —————–
def DateValidator1() -> DateValidator:
[…]
Column Disagreement Repaired show the number of disagreements from RQ1 (§4.2) which MatchFixAgent determined as not equivalent and successfully generated a correct patch. Of the $49$ disagreements resolved in favor of MatchFixAgent, we further asked our manual investigators if the generated patch was correct or not. On average, $\frac{47}{49}$ ($95.9\%$) of patches were validated as correct. Notice that this column cannot be directly compared with existing techniques, because they validated disagreements as functionally correct and did not generate any patches. Moreover, Column Coverage indicate the overall coverage for subject projects and the total improvement as a result of tests generated by MatchFixAgent. The generated tests help improve code coverage by $8.5\%$ (from $89.6\%$ to $98.1\%$). Project commons-fileupload from AlphaTrans sees the most improvement ($78.1\%$), with most of its translated fragments now validated with MatchFixAgent tests.
Summary. MatchFixAgent patches fix $50.6\%$ of translation bugs, $32.1\%$ more than existing repair techniques. Its generated target PL tests addresses the inadequate test suite limitation in prior work and help improve code coverage by $8.5\%$.
4.4. RQ3: Development Cost and Adaptability
In this RQ, we demonstrate the development cost of MatchFixAgent4.4.1) and its adaptability with other LLMs and agentic frameworks (§4.4.2).
Figure 5. Development cost of MatchFixAgent compared against existing tools.
4.4.1. Development Cost.
Figure 5 shows the development cost of MatchFixAgent against existing techniques. We define cost as the tool’s total lines of code (LoC). As illustrated in the figure, developing MatchFixAgent is cheap and the initial version only consists of 1,6501{,}650 LoC, supporting $6$ different PLs. Its dependence only on the Tree-Sitter (Tree-Sitter, 2025) parser makes it easy to support more languages using only $280$ LoC. By contrast, other tools, such as, Skel (Wang et al., 2025a), AlphaTrans (Ibrahimzada et al., 2025), and Oxidizer (Zhang et al., 2025) only support one PL pair and require significant engineering effort to adapt to more languages. More precisely, MatchFixAgent is $\times 2.3$, $\times 6.6$, and $\times 11.6$ cheaper than Skel, AlphaTrans, and Oxidizer, respectively. The static nature of MatchFixAgent makes it cost-effective and scalable, achieving better performance and revealing major limitations in existing tools.
4.4.2. Adaptability.
The architecture of MatchFixAgent is largely independent of any specific LLM or agent framework, making it easy to extend and integrate with more LLMs. In this research question, we investigate the extent to which replacing Anthropic’s Claude 3.7 Sonnet with OpenAI o4-mini-2025-04-16 (OpenAI, 2025), and Claude Code with Codex (Team, 2025k) agent, impacts the performance of MatchFixAgent. Due to the limited cost budget, we randomly sampled $96$ instances from all subject projects. While sampling, we controlled for equal contribution of equivalent and non-equivalent translations, resulting in $58$ and $38$ samples for each, respectively.
Figure 6 illustrates the results of this study. Our analysis indicates that MatchFixAgent with Claude Code and Codex agrees $73\%$ of the time against existing validation systems. Moreover, we also analyzed both agents’ behavior in terms of problem understanding and finding a solution. Our investigation of agent trajectories (footprint of agent actions) shows OpenAI’s Codex makes fewer actions to explore the codebase, and attempts to provide a decision faster. By contrast, Anthropic’s Claude Code agent first plans and reasons thoroughly about its tasks, specifically called TODO List by the agent, and then takes specific actions to perform each of its planned tasks.
Summary. MatchFixAgent is up to $\times 11.6$ times cheaper to develop than existing techniques. Also, it can be easily adapted to other LLMs and agent frameworks, e.g., OpenAI.
Figure 6. Agreement and dispute cases between tool validation system and MatchFixAgent with Claude and Codex agents.
4.5. RQ4: Ablation Study
We present two ablation studies to highlight the contribution of the semantic analyzer and the test generator agent in MatchFixAgent.
Figure 7. Impact of semantic analyzer and test generator agent in MatchFixAgent. A standalone baseline agent struggles validating translations as the projects grow in size.
4.5.1. Impact of Semantic Analyzer and Test Generator Agent.
To investigate the importance of semantic analyzer and test generator agent in MatchFixAgent, we evaluated a standalone baseline agent that uses the same LLM and agent framework, e.g., Claude Sonnet 3.7 (Team, 2025a) and Claude Code (Team, 2025c). In order to perform a controlled study, we created a sample from the original dataset where existing tools and MatchFixAgent verdicts agree with each other, meaning we collected all non-dispute instances. In total, 1,0911{,}091 instances, $862$ equivalent, and $229$ non-equivalent translations were selected.
Figure 7 shows the result of this ablation study. Accuracy indicates the ratio of baseline verdicts that agree with the existing tool and MatchFixAgent. On average, the validation accuracy drops by $42.3\%$, illustrating the importance of MatchFixAgent’s semantic analyzer and test generator components. Across all projects, the baseline agent only reproduced MatchFixAgent and tool results in the histogram project from Oxidizer (Zhang et al., 2025), achieving $100\%$ validation correctness. For the remaining projects, the agent’s accuracy dropped as low as $25\%$ in iceberg-java, which is the project with the largest number of lines of code. This study confirms that without semantic guidance and in-the-loop test generation, the LLM agents often fail to distinguish subtle differences and validate functional equivalence, especially when projects become larger and more restrictive languages, such as Rust are involved in the translation.
4.5.2. Impact of Semantic Analyzer.
We perform another study by removing only the semantic analysis results when prompting the test generator and verdict agents. We use the same set with 1,0911{,}091 instances from the previous ablation to perform this study. Figure 8 illustrates the result of our second ablation. The results indicate that the performance of MatchFixAgent significantly drops by $39.7\%$ without the six semantic analyses. Furthermore, we observe that the test generator agent without semantic analyzer is more costly and on average spends $3.9\%$ ($66.3K$ instead of $63.7K$), $6.2\%$ ($136K$ instead of $128K$), $4.2\%$ ($9.4M$ instead of $9.0M$), and $7.5\%$ ($61.28h$ instead of $56.71h$) more turns/interactions, input tokens, output tokens, and time, respectively. This is expected since the purpose of semantic analysis is to provide the necessary context for the test generator agent, without the need for the agent to explore and figure it out for itself. Since each semantic analysis is a single call to the LLM, it is much cheaper than the agents that iteratively analyze the codebase.
Summary. Removing the semantic analyzer and test generator significantly drops the effectiveness of MatchFixAgent by $42.3\%$. Moreover, removing the semantic analyzer alone decreases the accuracy by $39.7\%$, and increases the test generator cost, on average, by $5.4\%$.
Figure 8. Removing the semantic analyzer decreases the effectiveness of MatchFixAgent, while increasing token consumption, number of turns, and processing time.
5. Related Work
Translation Validation and Repair. Existing automated translation validation techniques either rely on test execution (Pan et al., 2024; Zhang et al., 2025; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Ziftci et al., 2025; Xue et al., 2025; Yang et al., 2024b), formal methods (Yang et al., 2024c; Nitin et al., 2024), or fuzzing (Eniser et al., 2024) for translation validation. Abid et al. (Abid et al., 2024) and AlphaTrans (Ibrahimzada et al., 2025) leverage GraalVM (Oracle, 2025) and language interoperability to execute code in both source and target PL for translation validation. Other tools like Oxidizer (Zhang et al., 2025) and Syzygy (Shetty et al., 2024) instrument programs to extract input-output pairs and use in target PL for validation. Skel (Wang et al., 2025a) validates translations through test execution by converting Python tests to JavaScript simply through string replacement. This suffices since source Python tests are simple function calls and value comparisons. For automated translation repair, most tools (Zhang et al., 2025; Wang et al., 2025a; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Yang et al., 2024b; Pan et al., 2024) rely on simple reprompting of LLMs with execution feedback, which has proven ineffective. Specifically, AlphaTrans (Ibrahimzada et al., 2025) performs independent reprompting of suspicious fragments based on execution trace, while Skel (Wang et al., 2025a) requires a user for manually repairing translation bugs.
LLM Agents. With the increasing prominence of agent-based frameworks (Liu et al., 2024; Xi et al., 2025), recent research and industrial efforts have turned towards leveraging these frameworks to address various software engineering tasks (Jimenez et al., 2024; Yang et al., 2025; Chowdhury et al., 2024). SWE-agent (Yang et al., 2024a) introduces a specialized agent-computer interface (ACI) facilitating agent interaction with code repositories via file reading, editing, and execution of bash commands. AutoCodeRover (Zhang et al., 2024), which provides LLM agents with specialized code-search APIs, enables iterative retrieval and localization of code segments associated with bugs. SpecRover (Ruan et al., 2025) enhances AutoCodeRover by emphasizing specification inference, generating function summaries, and providing targeted feedback at crucial agent execution stages. Agentless (Xia et al., 2025) further shows simple LLM agents can fix real-world bugs without using excessive tools and model complex environment behavior. Besides these state-of-the-art frameworks, numerous additional agent-based approaches exist both in open-source (Wei et al., 2025b; Ouyang et al., 2025; Bouzenia et al., 2024; Team, 2025h; AI, 2025) and commercial products (Wang et al., 2025b; Amazon, 2025; Cognition, 2025).
6. Threats to Validity
Similar to prior techniques, MatchFixAgent comes with some limitations and threats to the validity. In this section, we discuss how we mitigated various threats.
Internal Validity. There are two main threats to internal validity. First, we only run experiments once. Since LLMs are inherently non-deterministic, running experiments again may produce different results. While it is highly likely some individual equivalence verdicts and repair results would change if experiments were run again, it is highly unlikely the aggregate metrics we report would change significantly given the large number of translation samples we use (2,2192{,}219 pairs). Second, our human investigation does not assess ground truth equivalence. We only assess whether an inequivalent verdict was correct, but we do not analyze the correctness of equivalent verdicts. While this means we don’t have any measure of true accuracy of MatchFixAgent, we still can claim MatchFixAgent is more accurate than existing automated validation techniques.
External Validity. One main external threat is the generalizability of our approach. Our validation and repair system is very generic and can be extended to more PL pairs with minimal engineering effort. Also, the majority of tools that we used, for example, Tree-Sitter (Tree-Sitter, 2025) can support a large set of PLs. To mitigate external validity, we built the initial version of MatchFixAgent with six PLs.
Construct Validity. In order to minimize construct validity, MatchFixAgent is built on well-known and rigorously tested tools, e.g., Tree-Sitter, Claude Code, and Codex.
7. Conclusion
In this work, we presented MatchFixAgent, a language-agnostic neuro-symbolic technique which combines the power of program analysis and LLM agents for autonomous repository-level translation validation and repair. MatchFixAgent performs various semantic analyses—including control-flow and data-flow analyses—to systematically generate targeted tests, enabling demonstration of functional equivalence or detection of semantic bugs. Through rigorous evaluation on multiple benchmarks and different language-pairs, we show MatchFixAgent is effective in automatically validating and repairing translation pairs. It further generates high-quality reports that can be used by end-users for better understanding of translated programs and the validation process. Our manual investigation of generated reports reveals significant limitation of existing techniques in code translation validation and repair. MatchFixAgent is cost-effective and scalable, it only requires on average $280$ lines of code to support more PLs and it validates each instance in approximately $5$ minutes. To the best of our knowledge, MatchFixAgent is the first approach that can effectively validate and repair translations in repository-level across multiple PLs.

### v2

@@
 Program Analysis, Neuro-Symbolic Code Translation Validation and Repair, LLM Agent
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Foundations of Software Engineering; July 05–09, 2026; Montreal, Canadaisbn: 978-1-4503-XXXX-X/2018/06
 1. Introduction
@@
 The LLM is prompted also prompted to produce a plausible input that would trigger dissimilar IO behavior if it believes the translation is inequivalent.
-This methodology catches inequivalences such as differing error messages, inconsistent encoding assumptions, or missing side effects—often overlooked by structural analyses (§3.2.13.2.2) alone.
+This methodology catches inequivalences such as differing error messages, inconsistent encoding assumptions, or missing side effects—often overlooked by structural analyses (§3.2.1, §3.2.2) alone.
 3.2.4. Library Analyzer
@@
 textrank (Belicza, 2025) 52 1132 100 217 22
-AlphaTrans (Ibrahimzada et al., 2025) cli (Foundation, 2025b) Java Python 273 37841 100 372 201
-csv (Foundation, 2025c) 235 33072 100 392 278
-fileupload (Foundation, 2025d) 192 3567 100 246 185
-validator (Foundation, 2025e) 646 41605 95.5 216 164
+AlphaTrans (Ibrahimzada et al., 2025) cli (Foundation, 2025b) Java Python 273 37841 93.8 372 201
+csv (Foundation, 2025c) 235 33072 88.1 392 278
+fileupload (Foundation, 2025d) 192 3567 20.8 246 185
+validator (Foundation, 2025e) 646 41605 62.5 216 164
 Skel (Wang et al., 2025a) bst (Algorithms, 2025a) Python JavaScript 19 123 100 203000 47000
-colorsys (Team, 2025l) 8 120 96.3 67900 32300
-heapq (Team, 2025n) 22 189 78.1 67900 32300
-html (Team, 2025o) 44 684 65.3 67900 32300
-mathgen (Weiler, 2025) 81 735 93.8 711 183
-rbt (Algorithms, 2025b) 27 366 88.1 203000 47000
-strsim (Luo, 2025) 64 654 20.8 1014 125
-toml (Pearson, 2025) 72 1206 62.5 1126 192
+colorsys (Team, 2025l) 8 120 100 67900 32300
+heapq (Team, 2025n) 22 189 100 67900 32300
+html (Team, 2025o) 44 684 95.5 67900 32300
+mathgen (Weiler, 2025) 81 735 100 711 183
+rbt (Algorithms, 2025b) 27 366 96.3 203000 47000
+strsim (Luo, 2025) 64 654 78.1 1014 125
+toml (Pearson, 2025) 72 1206 65.3 1126 192
 RustRepoTrans (Ou et al., 2024) charset (Jawah, 2025) Python Rust 33 4231 100 672 56
@@
 The task for each benchmark problem is to give a verdict on the functional equivalence of pairs in two different PLs, and repair translation in case of equivalence.
-Our subjects are open-source repository-level translations with equivalence verdicts available [^4] We exclude rule-based transpilers as they do not propose a validation mechanism, i.e., their translation is (theoretically) correct by construction.
+Our subjects are open-source repository-level translations with equivalence verdicts available [^4] We exclude rule-based transpilers as they do not include a validation mechanism, i.e., their translation is (theoretically) correct by construction.
 from the peer-reviewed literature [^5] This criterion excludes techniques such as RustMap (Cai et al., 2025) and C2SaferRust (Nitin et al., 2025)..
@@
 The results indicate that the performance of MatchFixAgent significantly drops by $39.7\%$ without the six semantic analyses.
-Furthermore, we observe that the test generator agent without semantic analyzer is more costly and on average spends $3.9\%$ ($66.3K$ instead of $63.7K$), $6.2\%$ ($136K$ instead of $128K$), $4.2\%$ ($9.4M$ instead of $9.0M$), and $7.5\%$ ($61.28h$ instead of $56.71h$) more turns/interactions, input tokens, output tokens, and time, respectively.
+Furthermore, we observe that the test generator agent without semantic analyzer is more costly and on average spends $3.9\%$ ($66.3K$ instead of $63.7K$), $6.2\%$ ($136K$ instead of $128K$), $4.2\%$ ($9.4M$ instead of $9.0M$), and $7.5\%$ ($61.28\,\mathrm{h}$ instead of $56.71\,\mathrm{h}$) more turns/interactions, input tokens, output tokens, and time, respectively.
 This is expected since the purpose of semantic analysis is to provide the necessary context for the test generator agent, without the need for the agent to explore and figure it out for itself.
@@
 Translation Validation and Repair.
-Existing automated translation validation techniques either rely on test execution (Pan et al., 2024; Zhang et al., 2025; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Ziftci et al., 2025; Xue et al., 2025; Yang et al., 2024b), formal methods (Yang et al., 2024c; Nitin et al., 2024), or fuzzing (Eniser et al., 2024) for translation validation.
+Existing automated translation validation techniques either rely on test execution (Pan et al., 2024; Zhang et al., 2025; Ibrahimzada et al., 2025; Shetty et al., 2024; Ou et al., 2024; Ziftci et al., 2025; Xue et al., 2025; Yang et al., 2024b; Ke et al., 2025; Dehghan et al., 2025), formal methods (Yang et al., 2024c; Nitin et al., 2024), or fuzzing (Eniser et al., 2024) for translation validation.
 Abid et al.